OpenStudy (anonymous):

A street light is at the top of a 20 ft tall pole. A woman 6 ft tall walks away from the pole with a speed of 6 ft/sec along a straight path. How fast is the tip of her shadow moving when she is 45 ft from the base of the pole I get 36/14 and I redid it about 5 times and I keep getting the same, and I used related triangle... I'm out of things to do x.x

OpenStudy (anonymous):

Let x be the distance between the bottom of the pole and the woman. Let y be the length of her shadow. With a similar triangle you can get y/(x+y) = 6/20=3/10. Find y /x with the triangle Then Get the derivative for time, so dy/dt = (y/x) * dx/dt You know that dx/dt is 6ft/second, dy/dt is the relationship between (y/x) and dx/dt that i showed in the last step.